40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 472     Entry time: Fri May 9 08:40:24 2008
Author: steve 
Type: Update 
Category: SUS 
Subject: ETMY sus damping restored 
ETMY lost damping at 19:10 last night.
There was no seismic event than.
Sus damping was restored this morning.
ELOG V3.1.3-